# Moiré Superradiance in Cavity Quantum Electrodynamics with Quantum Atom Gas

### v2

@@ Abstract
The time evolution of $w(t)$ can be parametrized as $w(t)\sim t^{\nu}$ Larcher et al. (2009); Hiramoto and Abe (1988); Zhou et al. (2013). $\nu=1$ correspond to ballistic expansion. Under the combined effects of moiré lattice and cavity field quantum fluctuation, atoms would exhibit anomalous diffusion. We extract the time scaling $\nu$ from the atomic diffusion dynamics and present the results in Fig. 5(a). As expected, the value of $\nu$ decreases when pumping $\eta$ becomes stronger, signaling the system transfers from the regime of superdiffusion to that of subdiffusion. The boundary of the two regimes is identified with $\nu=\frac{1}{2}$. The superdiffusion to subdiffusion transfer is accompanied with the onset of superradiant phase transition and buildup of moiré lattice. This observation suggests a relation between atom diffusion and moiré pattern. As shown in Fig. 5(a), the scaling behavior indicate that superdiffusion to subdiffusion transition takes place at a smaller $\eta$ for the $M=5$ case (red-dashed line) as compared with that of $M=3$ (green-dashed line). This is consistent with our above analysis on the relation between moiré parameter and superradiant phase transition. Specifically, we examine two cases of M31 and M52, as indicated in Fig. 5(a), they subject to identical pumping of $\eta=0.46\eta_{c}$ however different $M$. M31 displays superdiffusion while M52 behaves subdiffusion, as demonstrated in Fig. 5(b). More detailed discussions on diffusion are given in the supplemental material sm .
@@
Discussion and outlook.— We have studied moiré effects on superradiant phase transition in a cold atom-cavity coupling system. This system resembles a generalized open Dicke model. The prominent role of moiré parameter in superradiant phase transition is explicitly explored. Implementing an optical lattice for 87Rb Bose-Einstein condensate inside a high-finesse optical cavity is feasible with existing technology Klinder et al. (2015b); Landig et al. (2016); Vaidya et al. (2018).
@@ Acknowledgements.
Supplementary Materials for
[…]
This PDF file includes:
Supplementary Notes I-VI
I. Matrix definition
II. Stationary equation solution
III. Excitations
IV. Cavity spectrum
V. Truncated Wigner methods
VI. Atom diffusion dynamics
Fig. S1
Fig. S2
I matrix definition
@@ VI. Atom diffusion dynamics
$\mathcal{K}=\left(\begin{array}[c]{ccccccc}0^{2}&&&&&&\\ &1^{2}&&&&&\\ &&2^{2}&&&&\\ &&&\ddots&&&\\ &&&&\ddots&&\\ &&&&&(n_{c}-1)^{2}&\\ &&&&&&n_{c}^{2}\\ &&&&&&\end{array}\right),~{}\mathcal{M}_{f_{n}}=\left(\begin{array}[]{ccccccc}&&&&\sqrt{2}&&\\ &&&1&&1&\\ &&\iddots&&&&\ddots\\ &1&&&&&\\ \sqrt{2}&&&&&&\\ &1&&&&&\\ &&\ddots&&&&\end{array}\right).$ (S1)
@@
II solve stationary equation
@@
$\dot{a}$ $=\left[i\left(\delta_{c}-\frac{U_{0}}{4}c^{\dagger}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})c\right)-\kappa\right]a-i\frac{\varepsilon_{p}}{2}c^{\dagger}\mathcal{M}_{f_{n+1}}c+a_{in},$
$\dot{c}$ $=-i\left[\omega_{k}\mathcal{K}+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}+\frac{U_{0}}{4}a^{\dagger}a(2\mathds{1}+\mathcal{M}_{2f_{n+1}})+\frac{\varepsilon_{p}}{2}(a+a^{\dagger})\mathcal{M}_{f_{n+1}}\right]c,$ (S2)
where the Gaussian noise operator $a_{in}$ has zero mean $\langle a_{in}(t)\rangle=0$ and nonvanishing correlation $\langle a_{in}(t)a_{in}^{\dagger}(t^{\prime})\rangle=\kappa\delta(t-t^{\prime})$. Inserting $a=\sqrt{N}\alpha+\delta a,~{}c_{0(n)}=e^{-i\mu t}\left[\sqrt{N}\psi_{0(n)}+\delta c_{0(n)}\right]$ into Eqs. (S2), the corresponding mean-field stationary equations can be derived as
$[\Delta(\psi)-i\kappa]\alpha+\frac{\eta}{2}\psi^{T}\mathcal{M}_{f_{n+1}}\psi=0,~{}\mathcal{C}(\alpha)\psi=\mu\psi,$ (S3)
@@
$\Delta(\psi)=-\Delta_{c}+u\psi^{T}\mathcal{M}_{2f_{n+1}}\psi,~{}\mathcal{C}(\alpha)=\omega_{k}\mathcal{K}+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}+u|\alpha|^{2}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})+\eta\operatorname{Re}(\alpha)\mathcal{M}_{f_{n+1}}.$ (S4)
Eqs. (S3) can be solved in a self-consistent manner. Start from an initial guess $\alpha$, diagonalize $\mathcal{C}(\alpha)$ to find out the $\psi$ corresponding to minimum $\mu$. Use this $\psi$ to derive a new $\alpha$, then repeat the above process until convergence is reached.
III excitations
Hiring linear pertubation theory, use the steady-state solutions $\left\{\alpha,\psi\right\}$ obtained from Section II, we first perform an orthogonal tranformation
$\delta b_{j}=\mathcal{O}^{T}\delta c_{j},$ (S5)
@@
$\frac{d}{dt}\delta a$ $=\left[-i\Delta\left(\psi\right)-\kappa\right]\delta a-i\sum_{j}\left(\frac{\eta}{2}g_{j}^{1}+u\alpha g_{j}^{2}\right)\left(\delta b_{j}+\delta b_{j}^{\dagger}\right)+a_{in},$
$\frac{d}{dt}\delta b_{j}$ $=-i\left(w_{j}-\mu\right)\delta b_{j}-i\left[\frac{\eta}{2}g_{j}^{1}\left(\delta a+\delta a^{\dagger}\right)+ug_{j}^{2}\left(\alpha^{\ast}\delta a+\alpha\delta a^{\dagger}\right)\right],$ (S6)
where $g_{j}^{1}=\mathcal{O}^{T}\mathcal{M}_{f_{n+1}}\psi$ and $g_{j}^{2}=\mathcal{O}^{T}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})\psi$. Since the quadrature $\frac{1}{2}\left(\delta b_{0}+\delta b_{0}^{\dagger}\right)$ is a constant of motion and the condensate phase fluctuation $\frac{1}{2i\sqrt{N}}\left(\delta b_{0}-\delta b_{0}^{\dagger}\right)$ would not take significant effect in the thermodynamic limit, we then neglect the fluctuation $\delta b_{0}$, deduce the Langevin equations for the operator vector $\mathcal{V}=\left\{\delta a,\delta a^{\dagger},\delta b_{1},\delta b_{1}^{\dagger},\dots,\delta b_{n_{c}},\delta b_{n_{c}}^{\dagger}\right\}^{T}$ as
$\frac{d}{dt}\mathcal{V}=\mathcal{M}_{e}\mathcal{V}+\xi,$ (S7)
@@
$\mathcal{M}_{e}=\left(\begin{array}[c]{ccccc}-i\Delta-\kappa&0&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&\cdots\\ 0&i\Delta-\kappa&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&\cdots\\ -i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&-i\left(w_{1}-\mu\right)&0&0\\ i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&0&i\left(w_{1}-\mu\right)&0\\ \vdots&\vdots&0&0&\ddots\end{array}\right),$ (S8)
and $\xi=\left(a_{in},a_{in}^{\dagger},0,0,\dots\right)^{T}$. A phase transition occurs if $\mathcal{M}_{e}$ inherits a zero energy eigenstate. We numerically diagonalize $\mathcal{M}_{e}$ to obtain a pair of lowest atom polariton excitations, with whose value approach 00 we determine the phase transition critical point, as shown in Fig. 2(b) of the main text.
In the normal phase $\alpha=0$, for small $V_{l}$ one can expect that the steady-state $\psi$ have most population in the homogeneous mode and a small excitation in the mode of $\cos(2k_{l}x)$, in the meanwhile the transformation matrix $\mathcal{O}$ is essentially identity. Then one can roughly only account $j=f_{n+1}$ and $2f_{n}-f_{n+1}$ in $g_{j}^{1}$. Physically they correspond to the process in which the atoms in the homogeneous mode and those in the mode of $\cos(2k_{l}x)$ are scattered by the cavity photon, respectively. In this case we can consider the following two block matrix
$\mathcal{M}_{ej}=\left(\begin{array}[c]{cccc}-i\Delta-\kappa&0&-i\frac{\eta}{2}g_{j}^{1}&-i\frac{\eta}{2}g_{j}^{1}\\ 0&i\Delta-\kappa&i\frac{\eta}{2}g_{j}^{1}&i\frac{\eta}{2}g_{j}^{1}\\ -i\frac{\eta}{2}g_{j}^{1}&-i\frac{\eta}{2}g_{j}^{1}&-i\left(w_{j}-\mu\right)&0\\ i\frac{\eta}{2}g_{j}^{1}&i\frac{\eta}{2}g_{j}^{1}&0&i\left(w_{j}-\mu\right)\end{array}\right),$ (S9)
@@
$\left(\begin{array}[c]{cccc}-i\Delta-\kappa&0&0&0\\ 0&i\Delta-\kappa&0&0\\ 0&0&-iw_{j}^{s}&0\\ 0&0&0&iw_{j}^{s}\end{array}\right)$ (S10)
@@
With the increase of the moiré parameter $M$, the value of $w_{j}-\mu$ will decrease as can be seen from the expression of $\mathcal{C}(\alpha)$ (S4) since $\omega_{k}=\omega_{r}/M^{2}$, rendering that $\eta_{c}$ also decreases. Physically this can be understood as an effective enlargement of the lattice constant, which will decrease the energy gap and thus facilitating phase transition.
IV cavity spectrum
@@
[…]
$=2\pi N\left|\alpha\right|^{2}\delta\left(\omega\right)+\left\langle\delta a^{\dagger}\left(\omega\right)\delta a\left(\omega^{\prime}\right)\right\rangle,$ (S11)
@@
We left multiply $\mathcal{O}^{l}$ on both sides of Eq. (S7), where the rows of $\mathcal{O}^{l}$ contain the left eigenvectors of the complex matrix $\mathcal{M}_{e}$, from which one can have
$\frac{d}{dt}\mathcal{V}_{l}=\mathcal{WV}_{l}+\xi_{l},$ (S12)
where $\mathcal{V}_{l}=\mathcal{O}^{l}\mathcal{V}=\left\{\delta d,\delta d^{\dagger},\delta e_{1},\delta e_{1}^{\dagger},\dots,\delta e_{n_{c}},\delta e_{n_{c}}^{\dagger}\right\}^{T}$, $\mathcal{W}$ is a diagonal matrix with the diagonal elements $\left\{-i\Delta-\kappa,i\Delta-\kappa,-iw_{1}^{s},iw_{1}^{s},\dots,-iw_{n_{c}}^{s},iw_{n_{c}}^{s}\right\}^{T}$, and $\xi_{l}^{j}=\mathcal{O}_{j,1}^{l}a_{in}+\mathcal{O}_{j,2}^{l}a_{in}^{\dagger}$. Introduce Fourier transforms to the quantum Langevin equations (S12) with
[…]
@@
$\delta e_{j}\left(\omega\right)$ $=\frac{\tilde{\xi}_{l}^{j+2}}{-i\left(\omega-w_{j}^{s}\right)+\gamma},\delta e_{j}^{\dagger}\left(-\omega\right)=\frac{\tilde{\xi}_{l}^{j+3}}{-i\left(\omega+w_{j}^{s}\right)+\gamma},$ (S14)
@@
$\delta a\left(\omega\right)=\mathcal{O}_{11}^{r}\delta d\left(\omega\right)+\mathcal{O}_{12}^{r}\delta d^{\dagger}\left(-\omega\right)+\sum_{j=1}^{n_{c}}\left[\mathcal{O}_{1,j+2}^{r}\delta e_{j}\left(\omega\right)+\mathcal{O}_{1,j+3}^{r}\delta e_{j}^{\dagger}\left(-\omega\right)\right].$ (S15)
Combining eq. (S14) and (S15), we can have
@@
$\simeq\sum_{j=1}^{n_{c}}\left[\frac{\kappa\delta\left(\omega-\omega^{\prime}\right)}{\left|-i\left(\omega-w_{j}^{s}\right)+\gamma\right|^{2}}\left|\mathcal{O}_{j+2,2}^{l}\mathcal{O}_{1,j+2}^{r}\right|^{2}+\frac{\kappa\delta\left(\omega-\omega^{\prime}\right)}{\left|-i\left(\omega+w_{j}^{s}\right)+\gamma\right|^{2}}\left|\mathcal{O}_{j+3,2}^{l}\mathcal{O}_{1,j+3}^{r}\right|^{2}\right],$ (S16)
@@
V truncated Wigner methods

### v3

@@ Abstract
For atomic wavepacket, the time evolution of its width $w(t)$ can be parametrized as $w(t)\sim t^{\nu}$ Larcher et al. (2009); Hiramoto and Abe (1988); Zhou et al. (2013). $\nu=1$ correspond to ballistic expansion. We extract the time scaling $\nu$ from the atomic diffusion dynamics and present the results in Fig. 7(a). The time scaling $\nu$ gradually decreases when pumping $\eta$ becomes stronger. This behavior is in contrast with the mean-field theory results, which predict an abrupt jump of $\nu$ from the value of $1$ (ballistic expansion) to approximately 0 (wavepacket localization) upon the onset of superradiant phase transition, signaling anomalous diffusion. Via mapping to the mean-field critical points, we identify two regimes of anomalous diffusion: superdiffusion ($\nu>\frac{1}{2}$) versus subdiffusion ($\nu<\frac{1}{2}$). The boundary separating these two regimes are determined by the intersection between the horizontal line of $\nu=\frac{1}{2}$ and the corresponding time scaling curves, as shown in Fig. 7(a). The superdiffusion to subdiffusion transition takes place at a smaller $\eta$ for the $M=5$ case (red-dashed line) as compared with that of $M=3$ (green-dashed line).
More specifically, we exemplify the diffusion dynamics in Fig. 7(b). With the increase of pumping $\eta$, corresponding to M51 ($M=5$, $\eta=0.4\eta_{c}$) $\rightarrow$ M52 ($M=5$, $\eta=0.46\eta_{c}$), and M31 ($M=3$, $\eta=0.46\eta_{c}$) $\rightarrow$ M32 ($M=3$, $\eta=0.5\eta_{c}$), one can observe that the diffusion property varies from superdiffusion to subdiffusion. Apart from that, moiré effect also matters in atomic diffusion. For M31 and M52 which subject to identical pumping of $\eta=0.46\eta_{c}$ however different $M$, M31 displays superdiffusion while M52 behaves subdiffusion, as demonstrated in Fig. 7(b). The anomalous diffusion behavior is amount to the combined effects of moiré lattice and cavity field quantum fluctuation.
VII discussion and outlook
@@
We have studied moiré effects on superradiant phase transition in a cold atom-cavity coupling system. This system resembles a generalized open Dicke model. The prominent role of moiré parameter in superradiant phase transition is explicitly explored. Implementing an optical lattice for 87Rb Bose-Einstein condensate inside a high-finesse optical cavity is feasible with existing technology Klinder et al. (2015b); Landig et al. (2016); Vaidya et al. (2018).
@@ Acknowledgements.
Appendix A Matrix definition
@@
$\mathcal{K}$ $=\left(\begin{array}[]{ccccccc}0^{2}&&&&&&\\ &1^{2}&&&&&\\ &&2^{2}&&&&\\ &&&\ddots&&&\\ &&&&\ddots&&\\ &&&&&(n_{c}-1)^{2}&\\ &&&&&&n_{c}^{2}\end{array}\right),$ (13)
$\mathcal{M}_{f_{n}}$ $=\left(\begin{array}[]{ccccccc}&&&&\sqrt{2}&&\\ &&&1&&1&\\ &&a&&&&\ddots\\ &1&&&&&\\ \sqrt{2}&&&&&&\\ &1&&&&&\\ &&\ddots&&&&\end{array}\right).$ (21)
@@
Appendix B Solve stationary equation
@@
$\dot{a}$ $=\left[i\left(\delta_{c}-\frac{U_{0}}{4}c^{\dagger}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})c\right)-\kappa\right]a$
$-i\frac{\varepsilon_{p}}{2}c^{\dagger}\mathcal{M}_{f_{n+1}}c+a_{in},$
$\dot{c}$ $=-i\left[\omega_{k}\mathcal{K}+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}+\frac{U_{0}}{4}a^{\dagger}a(2\mathds{1}+\mathcal{M}_{2f_{n+1}})\right$.
$\left.+\frac{\varepsilon_{p}}{2}(a+a^{\dagger})\mathcal{M}_{f_{n+1}}\right]c,$ (22)
where the Gaussian noise operator $a_{in}$ has zero mean $\langle a_{in}(t)\rangle=0$ and nonvanishing correlation $\langle a_{in}(t)a_{in}^{\dagger}(t^{\prime})\rangle=\kappa\delta(t-t^{\prime})$. Inserting $a=\sqrt{N}\alpha+\delta a,~c_{0(n)}=e^{-i\mu t}\left[\sqrt{N}\psi_{0(n)}+\delta c_{0(n)}\right]$ into Eqs. (22), the corresponding mean-field stationary equations can be derived as
$\Delta(\psi)-i\kappa]\alpha+\frac{\eta}{2}\psi^{T}\mathcal{M}_{f_{n+1}}\psi=0,~\mathcal{C}(\alpha)\psi=\mu\psi,$ (23)
@@
$\Delta(\psi)$ $=-\Delta_{c}+u\psi^{T}\mathcal{M}_{2f_{n+1}}\psi,$
$\mathcal{C}(\alpha)$ $=\omega_{k}\mathcal{K}+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}+u|\alpha|^{2}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})$
$+\eta\operatorname{Re}(\alpha)\mathcal{M}_{f_{n+1}}.$ (24)
Eqs. (23) can be solved in a self-consistent manner. Start from an initial guess $\alpha$, diagonalize $\mathcal{C}(\alpha)$ to find out the $\psi$ corresponding to minimum $\mu$. Use this $\psi$ to derive a new $\alpha$, then repeat the above process until convergence is reached.
Appendix C Excitations
Hiring linear pertubation theory, use the steady-state solutions $\left\{\alpha,\psi\right\}$ obtained from Section B, we first perform an orthogonal tranformation
$\delta b_{j}=\mathcal{O}^{T}\delta c_{j},$ (25)
@@ Appendix C Excitations
$\frac{d}{dt}\delta a$ $=\left[-i\Delta\left(\psi\right)-\kappa\right]\delta a-i\sum_{j}\left(\frac{\eta}{2}g_{j}^{1}+u\alpha g_{j}^{2}\right)$
$\times\left(\delta b_{j}+\delta b_{j}^{\dagger}\right)+a_{in},$
$\frac{d}{dt}\delta b_{j}$ $=-i\left(w_{j}-\mu\right)\delta b_{j}-i\left[\frac{\eta}{2}g_{j}^{1}\left(\delta a+\delta a^{\dagger}\right)\right$.
$\left.+ug_{j}^{2}\left(\alpha^{\ast}\delta a+\alpha\delta a^{\dagger}\right)\right],$ (26)
where $g_{j}^{1}=\mathcal{O}^{T}\mathcal{M}_{f_{n+1}}\psi$ and $g_{j}^{2}=\mathcal{O}^{T}(2\mathds{1}+\mathcal{M}_{2f_{n+1}})\psi$. Since $w_{0}=\mu$, rendering the quadrature $\frac{1}{2}\left(\delta b_{0}+\delta b_{0}^{\dagger}\right)$ a constant of motion, we then neglect the fluctuation $\delta b_{0}$, deduce the Langevin equations for the operator vector $\mathcal{V}=\left\{\delta a,\delta a^{\dagger},\delta b_{1},\delta b_{1}^{\dagger},\dots,\delta b_{n_{c}},\delta b_{n_{c}}^{\dagger}\right\}^{T}$ as
$\frac{d}{dt}\mathcal{V}=\mathcal{M}_{e}\mathcal{V}+\xi,$ (27)
@@
$\mathcal{M}_{e}=\left(\begin{array}[]{ccccc}-i\Delta-\kappa&0&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&\cdots\\ 0&i\Delta-\kappa&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&\cdots\\ -i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&-i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&-i\left(w_{1}-\mu\right)&0&0\\ i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha^{\ast}g_{1}^{2}\right)&i\left(\frac{\eta}{2}g_{1}^{1}+u\alpha g_{1}^{2}\right)&0&i\left(w_{1}-\mu\right)&0\\ \vdots&\vdots&0&0&\ddots\end{array}\right),$ (28)
and $\xi=\left(a_{in},a_{in}^{\dagger},0,0,\dots\right)^{T}$. We numerically diagonalize $\mathcal{M}_{e}$ to obtain pairs of atom polariton excitations $w_{j}^{s}$, with whose imaginary value approach 0 we determine the phase transition critical point, as shown in Fig. 2(b).
In the normal phase $\alpha=0$, for small $V_{l}$ one can expect that for the steady-state populations in the homogeneous mode and the mode of $\cos(2k_{l}x)$ are dominate, in the meanwhile the transformation matrix $\mathcal{O}$ is essentially identity. Then one can roughly only account $j=f_{n+1}$ and $2f_{n}-f_{n+1}$ in $g_{j}^{1}$. Physically they correspond to the process in which the atoms in the homogeneous mode and those in the mode of $\cos(2k_{l}x)$ are scattered by the cavity photon, respectively. In this case we can consider the following two block matrix
$\mathcal{M}_{ej}=\left(\begin{array}[]{cccc}-i\Delta-\kappa&0&-i\frac{\eta}{2}g_{j}^{1}&-i\frac{\eta}{2}g_{j}^{1}\\ 0&i\Delta-\kappa&i\frac{\eta}{2}g_{j}^{1}&i\frac{\eta}{2}g_{j}^{1}\\ -i\frac{\eta}{2}g_{j}^{1}&-i\frac{\eta}{2}g_{j}^{1}&-i\left(w_{j}-\mu\right)&0\\ i\frac{\eta}{2}g_{j}^{1}&i\frac{\eta}{2}g_{j}^{1}&0&i\left(w_{j}-\mu\right)\end{array}\right),$ (29)
@@
$\left(\begin{array}[]{cccc}-i\Delta-\kappa&0&0&0\\ 0&i\Delta-\kappa&0&0\\ 0&0&-iw_{j}^{s}&0\\ 0&0&0&iw_{j}^{s}\end{array}\right)$ (30)
@@
With the increase of the moiré parameter $M$, the value of $w_{j}-\mu$ will decrease as can be seen from the expression of $\mathcal{C}(\alpha)$ (24) since $\omega_{k}=\omega_{r}/M^{2}$, rendering that $\eta_{c}$ also decreases. Physically the behavior of critical pumping decreasing can be understood as an effective enlargement of the lattice constant, which will decrease the energy gap and thus facilitating phase transition.
Appendix D Cavity spectrum
@@
$S\left(\omega\right)$ $=2\pi N\left|\alpha\right|^{2}\delta\left(\omega\right)+\int dte^{-i\omega t}\left\langle\delta a^{\dagger}\left(t_{s}+t\right)\delta a\left(t_{s}\right)\right\rangle$
$=2\pi N\left|\alpha\right|^{2}\delta\left(\omega\right)+\left\langle\delta a^{\dagger}\left(\omega\right)\delta a\left(\omega^{\prime}\right)\right\rangle,$ (31)
@@
We left multiply $\mathcal{O}^{l}$ on both sides of Eq. (27), where the rows of $\mathcal{O}^{l}$ contain the left eigenvectors of the complex matrix $\mathcal{M}_{e}$, from which one can have
$\frac{d}{dt}\mathcal{V}_{l}=\mathcal{WV}_{l}+\xi_{l},$ (32)
where $\mathcal{V}_{l}=\mathcal{O}^{l}\mathcal{V}=\left\{\delta d,\delta d^{\dagger},\delta e_{1},\delta e_{1}^{\dagger},\dots,\delta e_{n_{c}},\delta e_{n_{c}}^{\dagger}\right\}^{T}$, $\mathcal{W}$ is a diagonal matrix with the diagonal elements $\left\{-i\Delta-\kappa,i\Delta-\kappa,-iw_{1}^{s},iw_{1}^{s},\dots,-iw_{n_{c}}^{s},iw_{n_{c}}^{s}\right\}^{T}$, and $\xi_{l}^{j}=\mathcal{O}_{j,1}^{l}a_{in}+\mathcal{O}_{j,2}^{l}a_{in}^{\dagger}$. Introduce Fourier transforms to the quantum Langevin equations (32) with
$\mathcal{O}\left(\omega\right)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{i\omega t}\mathcal{O}\left(t\right)dt,$
$\mathcal{O}^{\dagger}\left(-\omega\right)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{i\omega t}\mathcal{O}^{\dagger}\left(t\right)dt,$ (33)
@@
$\delta e_{j}\left(\omega\right)$ $=\frac{\tilde{\xi}_{l}^{j+2}}{-i\left(\omega-w_{j}^{s}\right)+\gamma},\delta e_{j}^{\dagger}\left(-\omega\right)=\frac{\tilde{\xi}_{l}^{j+3}}{-i\left(\omega+w_{j}^{s}\right)+\gamma},$ (34)
@@
$\delta a\left(\omega\right)$ $=\mathcal{O}_{11}^{r}\delta d\left(\omega\right)+\mathcal{O}_{12}^{r}\delta d^{\dagger}\left(-\omega\right)+\sum_{j=1}^{n_{c}}\left[\mathcal{O}_{1,j+2}^{r}\delta e_{j}\left(\omega\right)\right$.
$\left.+\mathcal{O}_{1,j+3}^{r}\delta e_{j}^{\dagger}\left(-\omega\right)\right].$ (35)
Combining eq. (34) and (35), we can have
@@
$\simeq\sum_{j=1}^{n_{c}}\left[\frac{\kappa\delta\left(\omega-\omega^{\prime}\right)}{\left|-i\left(\omega-w_{j}^{s}\right)+\gamma\right|^{2}}\left|\mathcal{O}_{j+2,2}^{l}\mathcal{O}_{1,j+2}^{r}\right|^{2}+\frac{\kappa\delta\left(\omega-\omega^{\prime}\right)}{\left|-i\left(\omega+w_{j}^{s}\right)+\gamma\right|^{2}}\left|\mathcal{O}_{j+3,2}^{l}\mathcal{O}_{1,j+3}^{r}\right|^{2}\right],$ (36)
@@
Appendix E Truncated Wigner methods